PT costan -1 sincot -1 x= root of x 2 +1/x 2 +2

PT costan-1 sincot-1x= root of x2+1/x2+2

Grade:12

2 Answers

noogler
8 years ago
cos(tan-1(sin(cot-1x)))=cos(tan-1(sin(sin-1(1/root of 1+x2))))                      cot-1x=tan-1(1/x)=sin-1(1/1+x2)1/2
                                =cos(tan-1(1/1+x2)1/2)                                            sin(sin-1x)=x
                               =cos(cos-1(1+x2/2+x2)1/2)                                        tan-1(1/1+x2)=cos-1(1+x2/2+x2)1/2
                               =(1+x2/2+x2)1/2
noogler
8 years ago
cos(tan-1(sin(cot-1x)))=cos(tan-1(sin(sin-1(1/root of 1+x2)))) 
                                  =cos(tan-1(1/1+x2)1/2)  
                                  =cos(cos-1(1+x2/2+x2)1/2
                                  =(1+x2/2+x2)½
 
 
cot-1x=tan-1(1/x)=sin-1(1/1+x2)1/2
 sin(sin-1x)=x
tan-1(1/1+x2)=cos-1(1+x2/2+x2)1/2